Shutdown epi2 SNF 2007-03-08 12:00:36: not mapping wafers correctly

maurice at snf.stanford.edu maurice at snf.stanford.edu
Tue Apr 17 17:28:28 PDT 2007


Robot arms straighted.
Blade replaced.
Susceptor ring replaced.
Chamber A dome replaced.
Wafers have been cycled in and out.
System baked.  Three Etch and Coats were run.
SiGe750-1 recipe was run.  No Haze.
5um-Pdope recipe was run.  Dep rate ~4um...as expected.
System up.




More information about the epi2-pcs mailing list